a father tells his son,"i was two years more than your present age when you were born.if the father is 40 years now,how old was his son 6 years ago
Answer by ankor@dixie-net.com(22740)

father tells his son,"i was two years more than your present age when you were born."
if the father is 40 years now,how old was his son 6 years ago.
:
let s = son's present age
father is 40 yrs old
"
then
"i was two years more than your present age when you were born."
40 - s = s + 2
40 - 2 = s + s
38 = 2s
s = 38/2
s = 19 yrs is son's present age
:
"old was his son 6 years ago."
19 - 6 = 13 yrs old